add (13/10r)+(11/4r)
Found 2 solutions by algebrahouse.com, rfer:
Answer by algebrahouse.com(1659)

3/10r + 11/4r
= 6/20r + 55/20r {got a common denominator, 20r, to add the fractions}
= 71/20r {combined numerators over common denominator}

For more help from me, visit: www.algebrahouse.com



Answer by rfer(16322)

26/20r+55/20r=81/20r=4 1/20r